OpenStudy (anonymous):

-tan^2(x)+sec^2(x)=1 Someone help me verify please!

OpenStudy (kainui):

Break this up into sines and cosines using the simplest identities you know. For example, \[\Large \sec(x)=\frac{1}{\cos(x)}\]

OpenStudy (anonymous):

Does tan=sin/cos?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So I would get -(sin(x))/cos(x)+1/cos(x)=1 ???

OpenStudy (larseighner):

Wow! This is a Pythagorean identity that is usually demonstrated geometrically. So I am waiting with bated breath for the answer.

OpenStudy (anonymous):

Simplified to give me -sinx+1/cosx, which is equivalent to cosx/cosx thanks to the 1st Pythagorean identity =1 !!!

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):